# Project page: https://github.com/zhang-tao-whu/PCM

### v3

@@ 4.1 Experiment Setup
Point Cloud Transformer [15] - - 93.2 - 2.9
Point-BERT [73] 83.1 - 93.2 - 22.1
Point-MAE [37] 85.2 - 93.8 - 22.1
@@
Detailed settings. We conduct experiments on four datasets: ScanObjectNN [52] classification, ModelNet40 [63] classification, ShapeNetPart [5] part segmentation, and S3DIS [1] semantic segmentation. We train PCM using the AdamW optimizer [33] with an initial learning rate of 1e-4, employing a Cosine Decay and a weight decay of 1e-4. We perform warmup for 5 epochs and use a batch size of 32, consistent across all tasks unless otherwise specified. We train PCM for 250 epochs on the ScanObjectNN and ModelNet40 datasets, for 300 epochs on ShapeNetPart, and for 100 epochs on S3DIS. For ScanObjectNN and ModelNet40, we followed the PointNeXt using 1024 points, randomly sampled during training, and using farthest point sampling during testing. For ShapeNetPart, 2,048 randomly sampled points with normals were used as input for training and testing. For S3DIS, 24,576 randomly sampled points were used as input for training. Following PointNeXt [41], PCM employs multi-step learning rate decay during training on ShapeNetPart, decaying at epochs 210 and 270, with a decay rate of 0.5. The experimental settings for PCM-Tiny are identical to PCM on all datasets. All ablation experiments are conducted using PCM as the default architecture, implemented on the ScanObjectNN dataset with training shortened to 125 epochs. Apart from this adjustment, all other settings are identical to the main experiment.
[…]
@@
PointNet [38] 83.7 80.4 3.6
CurveNet [64] 86.8 - -
ASSANet-L [40] 86.1 - -
Point Transformer [75] 86.6 83.7 7.8
PointMLP [35] 86.1 84.6 -
PointNet++ [39] 85.1 81.9 1.0
PointNeXt-S† [41] 86.7$\pm$0.0 84.4$\pm$0.2 1.0
PointNeXt-S (C=160) [41] 86.5$\pm$0.1 - 22.5
PointNeXt-S†(C=160) [41] 87.0$\pm$0.1 85.2$\pm$0.1 22.5
PointMamba [28] 86.0 84.4 17.4
@@
PCM†(ours) 87.1$\pm$0.2 85.6$\pm$0.1 40.6
@@
Table 4: 3D semantic segmentation in S3DIS. † indicates using DeLA [6] as the local feature extractor.
Method OA mAcc mIOU ceiling floor wall beam column window door table chair sofa bookcase board clutter
PointNet [38] - 49.0 41.1 88.8 97.3 69.8 0.1 3.9 46.3 10.8 59.0 52.6 5.9 40.3 26.4 33.2
PointCNN [27] 85.9 63.9 57.3 92.3 98.2 79.4 0.0 17.6 22.8 62.1 74.4 80.6 31.7 66.7 62.1 56.7
PointNeXt-S [41] 87.9 - 63.4 - - - - - - - - - - - - -
PointNeXt [41] 91.0 77.2 71.1 - - - - - - - - - - - - -
PTv1 [75] 90.8 76.5 70.4 - - - - - - - - - - - - -
PTv2 [60] 91.6 78.0 72.7 - - - - - - - - - - - - -
PTv3 [59] - - 74.7 - - - - - - - - - - - - -
DeLA [6] 92.2 80.1 74.1 96.0 98.7 87.7 0.0 55.2 61.5 76.0 84.3 93.3 85.3 79.4 81.2 65.2
PCM-Tiny (ours) 88.2 71.0 63.4 93.3 96.7 80.6 0.1 35.9 57.7 60.0 74.0 87.6 50.1 69.4 63.5 55.9
PCM-Tiny†(ours) 96.1 85.7 82.6 97.6 98.6 91.5 0.0 82.3 62.7 91.7 95.1 94.2 90.9 92.8 85.1 87.0
4.2 Main Results
@@
3D object classification in ScanObjectNN dataset. ScanObjectNet [52] is a challenging point cloud classification dataset containing 15,000 real scanned objects categorized into 13 classes. It is known for its noise and occlusion challenges. Following PointMLP [35] and PointNeXt [41], we conducted experiments on PB_T50_RS, the most challenging and commonly used ScanObjectNN variant. As shown in Tab. 2, PCM achieved an OA of 88.1 and a mAcc of 86.6 on ScanObjectNN, surpassing the SOTA method PointNeXt by 0.4 in OA and 0.8 in mAcc. Compared to PointMLP, Mamba layers demonstrated significantly stronger modeling capabilities than MLP, with PointNeXt surpassing PointMLP by 2.7 in OA and 2.7 in mAcc. By reducing the number of Mamba layers and channel size, PCM-Tiny achieved an OA of 86.9 and a mAcc of 85.0 with only 20% of the parameters of PCM. It is worth noting that PCM-Tiny, with only 52% of the parameters of PointMLP (6.9 M vs. 13.2 M), still outperformed PointMLP by 1.5 in OA and 1.1 in mAcc.
@@
3D object classification in ModelNet40 dataset. ModelNet40 [63] is a widely used synthetic 3D object classification dataset consisting of 40 categories, each with 100 unique CAD models. As shown in Tab. 2, PCM achieved an OA of 93.4 and a mAcc of 90.7, reaching a performance comparable to PointNeXt [41]. PCM-Tiny achieved an OA of 93.1 and a mAcc of 90.6 with approximately 20% of the parameters of PCM. However, due to the smaller scale and less challenging nature of ModelNet40, performance on this dataset is difficult to differentiate between different methods’ modeling capabilities significantly, with most methods’ OA concentrated between 93 and 94. We have reproduced the experiments of PointMLP [35] and obtained an OA of 93.6, indicating that the high accuracy of 94.1 requires multiple repetitions and selection of the best.
@@
3D object part segmentation in ShapeNetPart dataset. ShapeNetPart [5] is a widely used dataset for 3D object part segmentation. It comprises 16,880 models from 16 different shape categories and 50 part labels. Experimental results on the ShapeNetPart dataset are shown in Tab. 3. PCM achieves 87.0 Ins. mIoU and 85.3 Cls. mIoU without using any test augmentation strategies such as voting, surpassing PointNeXt [41] by 0.5 Ins. mIoU. With the use of voting, PCM achieves 87.1 Ins. mIoU and 85.6 Cls. mIoU, surpassing PointNeXt by 0.1 mIoU and 0.4 Cls. mIoU. Point-Tiny, with only 8.8 M parameters, achieves 86.9 Ins. mIoU and 85.0 Cls. mIoU, surpassing PointNeXt with 22.5 M parameters by 0.4 Ins. mIoU. PCM outperforms PointMLP [35] by 1.0 Ins. mIoU and 1.0 Cls. mIoU, demonstrating the significant potential of Mamba for 3D point cloud modeling.
@@
3D semantic segmentation in S3DIS dataset. S3DIS [1] is a large-scale indoor point cloud benchmark containing 6 large indoor areas, 271 rooms, and 13 semantic categories. We report evaluation results on area 5 in Tab. 4. PCM-Tiny achieved 88.2 OA, 71.0 mAcc, and 63.4 mIoU, showing a 0.3 OA improvement compared to PointNext-S and achieving comparable mIoU.
PCM can be combined with any point-based local modeling method to enhance local feature extraction capabilities without any architecture modifications. Specifically, substituting the original point features (such as colour, normals, etc.) with the hidden embeddings obtained from the point-based local modeling method for each point serves as PCM’s input. When using the more powerful local feature extractor DeLA [6], PCM-Tiny achieved 96.1 OA, 85.7 mAcc, and 82.6 mIoU, significantly surpassing the previous SOTA models DeLA [6] and PTv3 [59] by 8.5 mIoU and 7.9 mIoU, respectively. Compared to using DeLA [6] alone, the introduction of PCM brought effective global perception capabilities, resulting in substantial performance improvements across nearly all categories on the S3DIS dataset.
Table 5: Ablation studies on serialization strategies. Each mamba layer is assigned a serialization order and listed inside {}. "xyz", "xzy", "yxz", "yzx", "zxy", and "zyx" represent different variants of our proposed Consistent Traverse Serialization.
@@
4.3 Ablation Analysis and Visualization
@@
Serialization strategies. The key to applying Mamba for point cloud modeling is transforming point clouds into point sequences. As shown in Tab. 7, we conduct the ablation experiment with different serialization strategies. Similar performance is achieved when all Mamba layers’ inputs are serialized using a single order, whether it’s z-order, Hilbert-order, or our proposed xyz-order. However, significant performance gains are observed when more serialization strategies are employed. When the three variants of our proposed consistent traverse serialization, namely "xyz", "yzx", and "zxy", are used together, PCM demonstrates a performance improvement of 0.17 OA and 0.11 mAcc compared to using only the "xyz" variant. When all six variants of consistent traverse serialization are combined, PCM shows a performance improvement of 0.39 OA and 0.51 mAcc. When all six variants of consistent traverse serialization, as well as "Hilbert," "z," and "z-trans" serialization strategies, are combined, PCM achieves a performance improvement of 0.49 OA and 0.54 mAcc. Different serialization strategies allow different Mamba layers to observe point clouds from different perspectives, resulting in more robust modeling of point cloud features.
@@
Channel size. When the SSM-based method processes a token in the sequence, it relies solely on the hidden states and the input token, so the hidden states must have sufficient channel size to store global information. To investigate this, we conducted ablation studies on channel size, and the results are shown in Tab. 7. When the channel size is reduced from 768 to 384, PCM exhibits a performance decay of 1.12 OA and 1.41 mAcc. A significant performance decay of 2.56 OA and 3.76 mAcc is observed when the channel size is reduced to 96. We then attempted to reduce the channel size of early stages, and the results show that excessively reducing the channel size of early stages (from 768 to 96) still leads to a performance decay of 1.36 OA and 2.2 mAcc. However, moderately reducing the channel size of the first two stages (from 768 to 384) only results in minor performance decay but can save significant computation. Therefore, in our final configuration, the first two stages adopt a channel size of 384, while the last two stages use a channel size of 768.
@@
Positional embedding. We evaluate the impact of different positional encoding strategies, and the results are shown in Tab. 7. Initially, we experimented with rotary position embedding; however, it yielded the poorest performance. This is attributed to RoPE encoding solely the sequence order, unsuited for sparse and irregular point cloud data. Learnable positional embedding, a common practice in image sequence modeling, similarly encodes the sequence position and performs comparably to rotary positional encoding. Achieving favorable outcomes can be as straightforward as mapping point cloud spatial coordinates using a Linear layer as the positional encoding, resulting in a performance improvement of 0.37 OA and 0.7 mAcc compared to RoPE. Replacing the linear layer with a stronger MLP did not enhance performance. Nevertheless, the performance deteriorated due to overfitting when employing a separate Linear layer for each mamba layer.
Table 8: Ablation on numbers of order prompts.
@@
Table 9: Ablation on neighborhood points.
@@
Table 10: Ablation on points downsampling. The number of points at different stages is listed within {} and connected with -.
Points OA (%) mAcc (%)
{1024-1024-1024-1024} 87.35 85.71
{1024-512-256-128} 87.20 85.54
{512-256-128-128} 86.95 85.32
{512-256-128-64} 86.68 85.12
Figure 5: The visualization results of part segmentation on the ShapeNetPart dataset.
@@
Order prompts. To enhance the understanding of point cloud sequences by Mamba layers, we propose order prompts and conduct ablation experiments to validate their effectiveness, as shown in Tab. 10. When using only one order prompt, PCM demonstrates a performance improvement of 0.38 OA and 0.73 mAcc compared to not using any order prompt. Performance peaks when using six order prompts, resulting in a performance gain of 0.83 OA and 1.47 mAcc compared to no order prompt. However, further increasing the number of order prompts does not yield higher performance gains, although it still significantly outperforms not using any order prompt.
@@
Local features. 3D point cloud data are sparse and have low semantic density, making local features crucial for understanding point clouds. We conducted ablation experiments on computing local features using different numbers of neighboring points, and the results are shown in Tab. 10. When the number of neighboring points is set to 0, meaning no local features are computed, and only relying on Mamba layers to model the global features of the point cloud, PCM achieves an OA of 79.32 and a mAcc of 76.25. Using only four neighboring points to compute local features, PCM improves performance with an OA of 84.66 and a mAcc of 82.72, showing an increase of 5.34 OA and 6.47 mAcc compared to when local features are not used. PCM achieves the highest performance when using 12 points for computing local features. However, with a further increase in the number of neighboring points, performance decreases, indicating that the current local feature extraction mechanism, such as the geometric affine module in PointMLP, is not proficient at modeling the global features of point clouds. Therefore, combining local feature extraction modules and Mamba layers to model point clouds’ local and global features is a promising approach.
Point downsampling. Point cloud data exhibit significant redundancy; thus, appropriate point downsampling can substantially reduce computational costs with minimal loss in performance. We experimented with several downsampling schemes, and the results are shown in Tab. 10. Downsampling by a factor of 2 for all stages except the first stage resulted in only a slight performance decrease of 0.15 OA and 0.17 mAcc while significantly reducing computation. However, excessive downsampling, leaving only 64 points for the last stage, led to a performance drop of 0.67 OA and 0.59 mAcc. We ultimately adopted a downsampling strategy of line 2 for PCM.
5 Conclusion

### v4

@@ A Naive Mamba-based Point Cloud Network
PCT (Guo et al. 2021) - - 93.2 - 2.9
Point-BERT (Yu et al. 2022) 83.1 - 93.2 - 22.1
Point-MAE (Pang et al. 2022) 85.2 - 93.8 - 22.1
PTv3 (Wu et al. 2024) 86.4 83.9 - - -
@@
Table 2: 3D object classification in ScanObjectNN and ModelNet40. Averaged results in three random runs using $1024$ points as input without voting are reported.
We conduct experiments on four datasets: ScanObjectNN (Uy et al. 2019) and ModelNet40 (Wu et al. 2015) classification datasets, ShapeNetPart (Chang et al. 2015) part segmentation dataset, and S3DIS (Armeni et al. 2016) semantic segmentation dataset. For the detailed experiment settings, please refer to the supplementary materials.
@@
PointNet (Qi et al. 2017a) 83.7 80.4 3.6
CurveNet (Xiang et al. 2021) 86.8 - -
ASSANet-L (Qian et al. 2021) 86.1 - -
Point Transformer (Zhao et al. 2021) 86.6 83.7 7.8
PointMLP (Ma et al. 2022) 86.1 84.6 -
PointNet++ (Qi et al. 2017b) 85.1 81.9 1.0
PTv1 (Zhao et al. 2021) 86.6 83.7 -
PointNeXt-S (C=160) (Qian et al. 2022) 86.5$\pm$0.1 - 22.5
DeLA (Chen et al. 2023a) 87.0 85.8 7.5
SpoTr (Park et al. 2023) 87.2 85.4 -
PointMamba (Liang et al. 2024) 86.0 84.4 17.4
@@
Table 3: Part segmentation in ShapeNetPart.
Method OA mAcc mIOU
PointNet (Qi et al. 2017a) - 49.0 41.1
PointCNN (Li et al. 2018) 85.9 63.9 57.3
PointNeXt (Qian et al. 2022) 91.0 77.2 71.1
Strat. Trans. (Lai et al. 2022) 91.5 78.1 72.0
SPT (Robert, Raguet, and Landrieu 2023) - - 68.9
SpoTr (Park et al. 2023) 90.7 76.4 70.8
PTv1 (Zhao et al. 2021) 90.8 76.5 70.4
PTv2 (Wu et al. 2022) 91.6 78.0 72.7
DeLA (Chen et al. 2023a) 92.2 80.1 74.1
DeLA+X-3D (Sun et al. 2024) 92.2 80.1 74.3
KPConvX-L (Thomas et al. 2024) 91.7 78.7 73.5
PTv3 (Wu et al. 2024) - - 74.7
PTv3† (Wu et al. 2024) 91.4 78.4 72.3
PCM-Tiny (ours) 92.9 81.6 74.1
PCM-Tiny† (ours) 95.1 82.8 79.6
@@
Table 4: 3D semantic segmentation in S3DIS. † indicates using DeLA (Chen et al. 2023a) blocks as the additional local feature extractor.
Main Results
@@
3D object classification in ScanObjectNN dataset. ScanObjectNet (Uy et al. 2019) is a challenging point cloud classification dataset containing 15,000 real scanned objects categorized into 13 classes. It is known for its noise and occlusion challenges. Following PointMLP (Ma et al. 2022) and PointNeXt (Qian et al. 2022), we conducted experiments on PB_T50_RS, the most challenging and commonly used ScanObjectNN variant. As shown in Table 2, PCM achieved an OA of 88.1 and a mAcc of 86.6 on ScanObjectNN, surpassing the SOTA method PointNeXt and PTv3 by 0.4 and 1.7 in OA, respectively. Compared to PointMLP, Mamba layers demonstrated significantly stronger modeling capabilities than MLP, with PointNeXt surpassing PointMLP by 2.7 in OA and 2.7 in mAcc. By reducing the number of Mamba layers and channel size, PCM-Tiny achieved an OA of 86.9 and a mAcc of 85.0 with only 20% of the parameters of PCM. It is worth noting that PCM-Tiny, with only 52% of the parameters of PointMLP (6.9 M vs. 13.2 M), still outperformed PointMLP by 1.5 in OA and 1.1 in mAcc.
The superior performance of PCM compared to PointMLP demonstrates the importance of Mamba’s global modeling capability for point cloud analysis.
@@
3D object classification in ModelNet40 dataset. ModelNet40 (Wu et al. 2015) is a widely used synthetic 3D object classification dataset consisting of 40 categories, each with 100 unique CAD models. As shown in Table 2, PCM achieved an OA of 93.4 and a mAcc of 90.7, reaching a performance comparable to PointNeXt (Qian et al. 2022). PCM-Tiny achieved an OA of 93.1 and a mAcc of 90.6 with approximately 20% of the parameters of PCM. However, due to the smaller scale and less challenging nature of ModelNet40, performance on this dataset is difficult to differentiate between different methods’ modeling capabilities significantly, with most methods’ OA concentrated between 93 and 94. We have reproduced the experiments of PointMLP (Ma et al. 2022) and obtained an OA of 93.6, indicating that the high accuracy of 94.1 requires multiple repetitions and selection of the best.
@@
3D object part segmentation in ShapeNetPart dataset. ShapeNetPart (Chang et al. 2015) is a widely used dataset for 3D object part segmentation. It comprises 16,880 models from 16 different shape categories and 50 part labels. Experimental results on the ShapeNetPart dataset are shown in Table 3. PCM achieves 87.0 Ins. mIoU and 85.3 Cls. mIoU without using extra test augmentation strategies such as voting, surpassing PointNeXt (Qian et al. 2022) by 0.5 Ins. mIoU. Point-Tiny achieves 86.9 Ins. mIoU and 85.0 Cls. mIoU, surpassing PointNeXt by 0.4 Ins. mIoU. PCM outperforms PointMLP (Ma et al. 2022) by 1.0 Ins. mIoU and 1.0 Cls. mIoU, demonstrating the significant potential of Mamba for 3D point cloud modeling.
@@
3D semantic segmentation in S3DIS dataset. S3DIS (Armeni et al. 2016) is a large-scale indoor point cloud benchmark containing 6 large indoor areas, 271 rooms, and 13 semantic categories. PCM achieved 74.1 mIoU and 92.9 OA, surpassing PointNext (Qian et al. 2022) by 3.0 mIoU and 1.9 OA and exceeding PTv2 (Wu et al. 2022) by 1.4 mIoU and 1.3 OA. Moreover, PCM attained performance comparable to the current SOTA point-based method DeLA (Chen et al. 2023a) and transformer-based method PTv3 (Wu et al. 2024).
For better extraction of local point features, we cascade 4 DeLA blocks before PCM as an additional local feature extractor; please refer to the supplementary for details. When combining with the more powerful local feature extractor (Chen et al. 2023a), PCM-Tiny achieved 95.1 OA, 82.8 mAcc, and 79.6 mIoU, significantly surpassing the previous SOTA models DeLA (Chen et al. 2023a) and PTv3 (Wu et al. 2024) by 5.5 mIoU and 4.9 mIoU, respectively.
We also evaluate the performance of PTv3 with the same additional DeLA local feature extractor. However, the additional local feature extractor does not bring performance improvements to PTv3. This might be due to PTv3 performing attention within a window size of 1024, a limitation imposed by the quadratic computational complexity of transformers, which results in its limited global modeling capability.
@@
Table 5: Ablation studies on serialization strategies. Each mamba layer is assigned a serialization order and listed inside {}. "xyz", "xzy", "yxz", "yzx", "zxy", and "zyx" represent different variants of our proposed Consistent Traverse Serialization strategy.
@@
Ablation Analysis and Visualization
@@
Effect of serialization strategies. The key to applying Mamba for point cloud modeling is transforming point clouds into point sequences. As shown in Table 5, we conduct the ablation experiment with different serialization strategies. Similar performance is achieved when all Mamba layers’ inputs are serialized using a single order, whether it’s z-order, Hilbert-order, or our proposed xyz-order. However, significant performance gains are observed when more serialization strategies are employed. When the three variants of our proposed consistent traverse serialization, namely "xyz", "yzx", and "zxy", are used together, PCM demonstrates a performance improvement of 0.17 OA and 0.11 mAcc compared to using only the "xyz" variant. When all six variants of consistent traverse serialization are combined, PCM shows a performance improvement of 0.39 OA and 0.51 mAcc. When all six variants of consistent traverse serialization, as well as "Hilbert," "z," and "z-trans" serialization strategies, are combined, PCM achieves a performance improvement of 0.49 OA and 0.54 mAcc. Different serialization strategies allow different Mamba layers to observe point clouds from different perspectives, resulting in more robust modeling of point cloud features.
@@
Impact of channel size. When the SSM-based method processes a token in the sequence, it relies solely on the hidden states and the input token, so the hidden states must have sufficient channel size to store global information. To investigate this, we conducted ablation studies on channel size, and the results are shown in Table 6. When the channel size is reduced from 768 to 384, PCM exhibits a performance decay of 1.12 OA and 1.41 mAcc. A significant performance decay of 2.56 OA and 3.76 mAcc is observed when the channel size is reduced to 96. We then attempted to reduce the channel size of early stages, and the results show that excessively reducing the channel size of early stages (from 768 to 96) still leads to a performance decay of 1.36 OA and 2.2 mAcc. However, moderately reducing the channel size of the first two stages (from 768 to 384) only results in minor performance decay but can save significant computation. Therefore, in our final configuration, the first two stages adopt a channel size of 384, while the last two stages use a channel size of 768.
@@
Ablation on positional embedding. We evaluate the impact of different positional encoding strategies, and the results are shown in Table 7. Initially, we experimented with rotary position embedding; however, it yielded the poorest performance. This is attributed to RoPE encoding solely the sequence order, unsuited for sparse and irregular point cloud data. Learnable positional embedding, a common practice in image sequence modeling, similarly encodes the sequence position and performs comparably to rotary positional encoding. Achieving favorable outcomes can be as straightforward as mapping point cloud spatial coordinates using a Linear layer as the positional encoding, resulting in a performance improvement of 0.37 OA and 0.7 mAcc compared to RoPE. Replacing the linear layer with a stronger MLP did not enhance performance. Nevertheless, the performance deteriorated due to overfitting when employing a separate Linear layer for each mamba layer.
@@
Table 8: Ablation on numbers of order prompts.
@@
Table 9: Ablation on neighborhood points.
@@
Order prompts. To enhance the understanding of point cloud sequences by Mamba layers, we propose order prompts and conduct ablation experiments to validate their effectiveness, as shown in Table 9. When using only one order prompt, PCM demonstrates a performance improvement of 0.38 OA and 0.73 mAcc compared to not using any order prompt. Performance peaks when using six order prompts, resulting in a performance gain of 0.83 OA and 1.47 mAcc compared to no order prompt. However, further increasing the number of order prompts does not yield higher performance gains, although it still significantly outperforms not using any order prompt.
@@
Local features. 3D point cloud data are sparse and have low semantic density, making local features crucial for understanding point clouds. We conducted ablation experiments on computing local features using different numbers of neighboring points, and the results are shown in Table 9. When the number of neighboring points is set to 0, meaning no local features are computed, and only relying on Mamba layers to model the global features of the point cloud, PCM achieves an OA of 79.32 and a mAcc of 76.25. Using only four neighboring points to compute local features, PCM improves performance with an OA of 84.66 and a mAcc of 82.72, showing an increase of 5.34 OA and 6.47 mAcc compared to when local features are not used. PCM achieves the highest performance when using 12 points for computing local features. However, with a further increase in the number of neighboring points, performance decreases, indicating that the current local feature extraction mechanism, such as the geometric affine module in PointMLP, is not proficient at modeling the global features of point clouds. Therefore, combining local feature extraction modules and Mamba layers to model point clouds’ local and global features is a promising approach.
Conclusion